Estimate the price impact of a 50-basis point interest rate change using a linear approximation of a 10-year 7% semi-annual coupon bond priced at 97.50.

a. 1.72

b. 3.40

c. 3.44

d. 6.80

e. 6.88
Answer #1

The bond prices have inverse relation with change in interest rate , i.e. the price moves in the opposite direction of the change in interest rates.

Value of bond = Maturity Couponi

Face Value $                100
Annual Coupon Rate 7.000%(annual) 7%/2=3.5%(semi annual)
Years to Maturity 10.00
Price of bond                   97.5
Payment Frequency                       2
YTM 7.36% using rate() function
Rate Price of Bond
7.357% $97.50
7.857% $94.14
7% $100.00
7.50% $96.53
8% $93.20
8.357% $90.92

after running the regression , output for the linear relation establish as

Price of bond = A + B* change in interest

Coefficients Standard Error t Stat P-value Lower 95%
Intercept 146.7565964 0.481033181 305.0862 6.93E-10 145.421
Interest rate Variable 1 -669.075076 6.254155755 -106.981 4.58E-08 -686.439

for 50 bps = 0.5%

Change = -669.07 * 0.5% = -3.34

Option C : 3.44